# Topological Obstructions to Dynamical Convexity

### v1

@@ 3. Symplectic Homology
In this section we recall basic facts about symplectic homology. We refer to the the following surveys and papers for a nice introduction with more details, [cieliebak2018symplectic],[lazarev2020contact],[oancea2004survey],[seidel2008biased]. We follow sign and grading conventions as in [lazarev2020contact].
@@
For our choice of $H\in H_{adm}$, the periodic trajectories $\mathcal{P}(H)$of the Hamiltonian vector field $X_{H}$ correspond to Morse critical points on $W$ and Reeb orbits on a level $r$ in $\hat{W}/W$ of period $s$ since $X_{H}=sR_{\alpha}.$ The Hamiltonian orbits are degenerate at the cylindrical end and we consider a small time-dependent pertubation of $H$, so that each $S^{1}$ family degenerates into two orbits[bourgeois2009exact].
@@
The symplectic homology groups or positive symplectic homology groups do not depend on the contact form of $\partial W$ (cf [ritter2010deformations, section 3.9] ). More precisely, for a contact form $\alpha$ on $\partial W$, one can show that W{W} is Liouville isomorphic to a Liouville domain $W^{\prime}$ which induces the contact form $\alpha$ on $\partial W^{\prime}$ upto a strict contactomorphism.
@@
The positive symplectic homology groups although being generated by Reeb orbits are not in general invariants of the contact boundary. It depends on the filling as the Floer trajectories can enter the filling. However it can be shown that for contact manifolds that admit dynamically convex or more generally ADCADC-contact forms as in [zhou2020vanishing], the positive symplectic homology is independent of fillings $W$ such that $\pi_{1}(\partial W)\rightarrow\pi_{1}(W)$ is injective and $c_{1}(W,\mathbb{Q})$ is torsion.
@@
Note that for a Liouville domain $W$ with contact boundary $(\partial W,\xi)$, $c_{1}(\xi,\mathbb{Q})=i^{*}c_{1}(TW,\mathbb{Q})$. Therefore $c_{1}(TW,\mathbb{Q})=0$ would imply $c_{1}(\xi,\mathbb{Q})=0$. For Weinstein domains of dimension greater than $6$, the converse is also true and we replicate the argument in [lazarev2020contact] with rational coefficients.
@@
Note that the vanishing of positive symplectic homology is an invariant of dynamically convex contact manifolds when the manifold is simply connected. This is a consequence of the fact that a manifold that is simply connected and dynamically convex is by definition strongly ADC and therefore linearized contact homology is well defined for the contact manifold (it has a unique augmentation) [chaidez2024contact], which vanishes if and only if the positive symplectic homology vanishes for fillable manifolds due to the existence of the Gysin sequence after identifying linearized contact homology with positive $S^{1}$-equivariant symplectic homology.
Proposition 3.11 ([viterbo1999functors]).
@@
The above maps are used to show that attaching a subcritical handle induces an isomorphism in $SH$. For more details see [fauck2016handle] or [cieliebak2002handle].
For cotangent bundles $T^{*}M$, Viterbo showed that there is an isomorphism between the $SH(T^{*}M)$ and the homology of the loop space $H_{*}(\mathcal{L}M)$ with $\mathbb{Z}/2\mathbb{Z}$ coefficients. In fact this isomorphism can be lifted to $\mathbb{Z}$ coefficients after twisting by a suitable local system([abouzaid2013symplectic]). However, the isomorphism with $\mathbb{Z}/2\mathbb{Z}$ suffices for us. Note that for a cotangent bundle $T^{*}M$ of a manifold $M$, we have $c_{1}(T^{*}M)=0$ and conventions are chosen so that the gradings match.
Theorem 3.12 ([viterbo1999functors]).
@@ 4. Unit Cotangent bundles and dynamical convexity
If the unit cotangent bundle $ST^{*}M$ of a simply connected manifold $M$ admits a non-degenerate dynamically convex contact form $\lambda$, then the map $\iota_{*}:H_{k}(M;\mathbb{Z}_{2})\rightarrow H_{k}(\mathcal{L}M;\mathbb{Z}_{2})$ induced by the inclusion of constant loops, is an isomorphism for $k\leq n-1$.
@@
Theorem 4.6 ([serre1951homologie][leray1950anneau]).
@@
Corollary 4.7 ([mccleary2001user]Example 55D).
@@
Since the edge homomorphism $H_{2}(\mathcal{L}M)\rightarrow E^{\infty}_{(2,0)}\rightarrow\cdots\rightarrow E^{2}_{(2,0)}=H_{2}(M)$ in the spectral sequence for the loop-loop fibration coincides with the fibration map $\pi_{*}$ [miller2021lectures, Proposition 64.2] , we have that $E^{2}_{(2,0)}=H_{2}(M)$ persists till the limiting page. This implies that $d_{2}=0$. The fact that the differential $d_{2}$ vanishes forces $E^{2}_{(0,1)}=H_{1}(\Omega M)$ to persist until the last page. The isomorphism $H_{2}(\mathcal{L}M)=H_{2}(M)$ implies that in the limiting page $H_{2}(M)$ is the sole contributor to $H_{2}(\mathcal{L}M)$. Therefore, $H_{1}(\Omega M)=0$.
@@
Theorem 4.10 (Theorem 1.3).
The unit cotangent bundle $ST^{*}M$ of a simply connected closed manifold $M$ (not necessarily orientable) with standard contact structure cannot be dynamically convex.
@@
Theorem 4.1 implies $M$ is simply connected and since strongly dynamically convex manifolds are dynamically convex, we obtain the result, using theorem 4.10. ∎
@@ 5. Dynamical convexity and surgery
The idea to obstruct dynamical convexity after surgery on Weinstein fillable contact manifolds is to observe the changes in $SH^{+}$ and in the homology after performing surgery. We refer to [fauck2020manifolds] for details on how attaching a handle induces maps on the canonical long exact sequence of symplectic homology. In particular, the author shows that, after handle attachment, the Viterbo transfer map 3.11 still respects the action filtration and therefore induces a map on the long exact sequence 3.5.
Note that $SH=\bigoplus_{\gamma\in\pi_{1}}SH_{\gamma}$ where $SH_{\gamma}$ denotes the subgroup generated by orbits in the homotopy class of $\gamma$. In this section we only consider the subgroup generated by contractible orbits and denote it by $SH$ or $SH^{+}$ for the sake of brevity. We also only consider $SH,SH^{+}$ and homology/cohomology groups over $\mathbb{Q}$ coefficients but omit the coefficients from the notation.
@@
The following proofs in this also hold for non-Weinstein fillable contact manifolds if the fillings have vanishing rational first chern class.
@@
The maps induced by surgery on the canonical long exact sequence [fauck2020manifolds, section 2.8] are
@@
we note that, since attaching a 1-handle is subcritical, the second and last arrows on symplectic homology are isomorphisms. Also note that since connectedness is preserved $H^{0}$ remains invariant and therefore $\operatorname{rk}(H^{1}(W^{\prime}))=\operatorname{rk}(H^{1}(W))+1$. The map $H^{1}(W^{\prime})\rightarrow H^{1}(W)$, induced by inclusion, is surjective by the long exact sequence in cohomology in section 5. Therefore, by the 4-lemma ([maclane2012homology, Chapter 12 Lemma 3.1]), the map $SH^{+}_{n}(W^{\prime})\rightarrow SH^{+}_{n}(W)$ is a surjection. The vanishing of $SH^{+}_{n}(W^{\prime})$ implies that $SH^{+}_{n}(W)$ vanishes.
@@
If $M$ is simply connected, dynamically convex and of dimension at least $5$, then by [kwon2024dynamically] $M$ is diffeomorphic to a sphere if the filling has vanishing $SH$. This follows from the h-cobordism theorem. Note that vanishing of $SH^{+}$ is independent of the filling by 3.10. ∎
@@
We can consider the process of connected sum as attaching a one handle to two components of $M$, thus decreasing the rank of $H^{0}$ by 1 and preserving the ranks of all the higher cohomology groups. $SH$ remains invariant as we are performing subcritical surgery.
Now, if $SH^{+}_{k}(W)=0$ for $k<n+1$, then $SH^{+}_{k}(DT^{*}N)=0$ for $k<n+1$. The proof of lemma 4.2 holds and this would imply $H_{k}(N,\mathbb{Z}/2\mathbb{Z})\hookrightarrow H_{k}(\mathcal{L}N,\mathbb{Z}/2\mathbb{Z})$ under inclusion is an isomorphism for $k<n+1$. Since $N$ is simply connected the argument following Lemma 4.2 in section 4 using the Serre spectral sequence would give a contradiction. Therefore there exists $k<n+1$ such that $SH^{+}_{k}(W)=0$.

### v2

@@ 3. Symplectic Homology
In this section we recall basic facts about symplectic homology before using symplectic homology of covering spaces to prove theorem 1.18. We refer to the the following surveys and papers for a nice introduction with more details, [CO18],[LAZ20],[OAN04],[SEI08]. We follow sign and grading conventions as in [LAZ20], [CO18].
@@
For our choice of $H\in H_{adm}$, the periodic trajectories $\mathcal{P}(H)$of the Hamiltonian vector field $X_{H}$ correspond to Morse critical points on $W$ and Reeb orbits on a level $r$ in $\hat{W}/W$ of period $s$ since $X_{H}=sR_{\alpha}.$ The Hamiltonian orbits are degenerate at the cylindrical end and we consider a small time-dependent pertubation of $H$, so that each $S^{1}$ family degenerates into two orbits[BO09].
@@
The symplectic homology groups or positive symplectic homology groups do not depend on the contact form of $\partial W$ (cf [RIT10, section 3.9] ). More precisely, for a contact form $\alpha$ on $\partial W$, one can show that W{W} is Liouville isomorphic to a Liouville domain $W^{\prime}$ which induces the contact form $\alpha$ on $\partial W^{\prime}$ upto a strict contactomorphism.
@@
The positive symplectic homology groups although being generated by Reeb orbits are not in general invariants of the contact boundary. It depends on the filling as the Floer trajectories can enter the filling. However it can be shown that for contact manifolds that admit dynamically convex or more generally ADCADC-contact forms as in [ZHO20], the positive symplectic homology is independent of fillings $W$ such that $\pi_{1}(\partial W)\rightarrow\pi_{1}(W)$ is injective and $c_{1}(W,\mathbb{Q})$ is torsion. These fillings are called topologically simple.
@@
Note that for a Liouville domain $W$ with contact boundary $(\partial W,\xi)$, $c_{1}(\xi,\mathbb{Q})=i^{*}c_{1}(TW,\mathbb{Q})$. Therefore $c_{1}(TW,\mathbb{Q})=0$ would imply $c_{1}(\xi,\mathbb{Q})=0$. For Weinstein domains of dimension greater than $6$, the converse is also true and we replicate the argument in [LAZ20] with rational coefficients.
@@
Note that the vanishing of positive symplectic homology is an invariant of dynamically convex contact manifolds when the manifold is simply connected. This is a consequence of the fact that a manifold that is simply connected and dynamically convex is by definition strongly ADC and therefore linearized contact homology is well defined for the contact manifold (it has a unique augmentation) [CHA24], which vanishes if and only if the positive symplectic homology vanishes for fillable manifolds due to the existence of the Gysin sequence after identifying linearized contact homology with positive $S^{1}$-equivariant symplectic homology.
Proposition 3.11 ([VIT99]).
@@
The above maps are used to show that attaching a subcritical handle induces an isomorphism in $SH$. For more details see [FAU16] or [CIE02].
For cotangent bundles $T^{*}M$, Viterbo showed that there is an isomorphism between the $SH(T^{*}M)$ and the homology of the loop space $H_{*}(\mathcal{L}M)$ with $\mathbb{Z}/2\mathbb{Z}$ coefficients. In fact this isomorphism can be lifted to $\mathbb{Z}$ coefficients after twisting by a suitable local system([ABO13]). However, the isomorphism with $\mathbb{Z}/2\mathbb{Z}$ suffices for us. Note that for a cotangent bundle $T^{*}M$ of a manifold $M$, we have $c_{1}(T^{*}M)=0$ and conventions are chosen so that the gradings match.
Theorem 3.12 ([VIT99]).
@@
3.1. Symplectic homology of covering spaces.
Zhou [ZHO21, Section 3.3][ZHO23] defines and uses symplectic homology of covering spaces to study fillings of asymptotically dynamically convex contact manifolds. Note from here onward till the end of the section, we use symplectic homology generated with contractible Reeb orbits. Given a covering $\tilde{W}\xrightarrow{\pi}W$, SH(W~)SH($\tilde${W}) is defined as the group generated by lifts of the periodic orbits of $W$ and the differential is determined by counting the moduli spaces of (unique) lifted trajectories. More precisely the chain group is defined as $SC_{*}(\tilde{W}):=\bigoplus_{x}\Pi_{\pi(\tilde{x})=x}\tilde{x}$ and it could be an infinite product in the fiber direction. The canonical map $x\rightarrow\Pi_{\pi(\tilde{x})=x}\tilde{x}$ induces maps on the level of $SH^{<\epsilon},SH,SH^{+}$ and fits into the long exact sequence of symplectic homology [ZHO23, 4.1]. Furthermore the map is a unital ring map in $SH$.
$H^{*}(W)$$SH_{*-n}(W)$$SH^{+}_{*-n}(W)$${H^{*}(\tilde{W})}$${SH_{*-n}(\tilde{W})}$${SH^{+}_{*-n}(\tilde{W})}$}{{...}}
Remark 3.15.
Zhou[ZHO21, Section 3.3][ZHO23] uses cohomological convention and defines symplectic cohomology groups. They coincide with our definition if we define our grading using $\mu_{CZ}$ instead of $\mu_{CZ}$
Zhou [ZHO23, Remark 4.1] also discusses symplectic homology twisted by $\mathbb{Z}[\pi_{1}]$ coefficients and makes a comparison with the symplectic homology of covering spaces. In this case, the chain group is the $\mathbb{Z}[\pi_{1}]$ module generated by the periodic orbits with a twisted differential by the holonomy action of the fundamental group. When $\pi_{1}$ is finite, symplectic homology of the universal covering space coincides with the symplectic homology twisted with $\mathbb{Z}[\pi_{1}]$ coefficients since the covering space is compact and SH(W~)SH($\tilde${W}) is generated by finitely many orbits. When $\pi_{1}$ is infinite, the groups are different since symplectic homology of the universal covering space has infinitely many generators in a fiber direction.
Remark 3.16.
[…]
Theorem 3.17.
Given a Liouville domain $W$, the pair of pants product makes the symplectic homology of $W$, twisted by $\mathbb{Z}[\pi_{1}]$ local coefficients a module over SH(W~)SH($\tilde${W}), the symplectic homology of the universal covering $\tilde{W}$.
Proof.
We notice that the chain complex of symplectic homology with $\mathbb{Z}[\pi_{1}]$ local coefficients is a subcomplex of that of symplectic homology of the universal cover SH(W~)SH($\tilde${W}) since its differential coincides with that of $SH(W,\mathbb{Z}[\pi_{1}])$ when restricted to finitely many orbits. Now let $z$ be the pair of pants product with an element $y$ represented by an infinite sum of orbits in the fiber direction and a single orbit $x$ . The projection to $W$ of the all the possible pair of pants between $x$ and summands in $y$ gives finitely many pair of pants in $W$ who have finite lifts after fixing $x$ due to unique lifting property as we are considering contractible orbits. Therefore $z\in SC(W,\mathbb{Z}[\pi_{1}])$ and we have a module action $SH(\tilde{W})\otimes SH(W,\mathbb{Z}[\pi_{1}])\rightarrow SH(W,\mathbb{Z}[\pi_{1}])$. ∎
Corollary 3.18.
Given a Liouville domain $W$, $SH(W)=0$ implies $SH(W,\mathbb{Z[\pi_{1}]}=0.$
Proof.
[…]
Theorem 3.19.
1.18 Let $M$ be a dynamically convex contact manifold such that there exists a topologically simple filling $W^{\prime}$ of $M$ with $SH(W^{\prime})=0$. Then for any topologically simple filling $W^{\prime}$ of $M$, $\pi_{k}(W)\cong\pi_{k}(M)$ for 1<k<2n11<k<2n-1 if $\pi_{1}(\partial W)\rightarrow\pi_{1}(W)$ is surjective.
Proof.
We remark that by [ZHO20], the vanishing of symplectic homology for any topologically simple filling implies the vanishing of symplectic homology for all topologically simple fillings. Therefore, $SH(W)=0$. Let $\tilde{W}$ denote the universal covering of $W$. When $\pi_{1}(\partial W)\rightarrow\pi_{1}(W)$ is surjective, the boundary $\partial\tilde{W}$ is connected. Examining the long exact sequence of the relative homotopy groups of the pair (W~,W~)($\partial\tilde$$\tilde{W}$), $\rightarrow\pi_{1}(\tilde{W})\rightarrow\pi_{1}(\tilde{W},\partial\tilde{W})\rightarrow\pi_{0}(\partial\tilde{W})\rightarrow\pi_{0}(\partial W)$, we deduce that $\pi_{1}(\tilde{W},\partial\tilde{W})=0.$
[…]
${H_{*+n}(\tilde{W},\partial\tilde{W})}$$SH_{*}(W,\mathbb{Z}[\pi_{1}])$$SH_{*}^{+}(W,\mathbb{Z}[\pi_{1}])$${H_{*+n-1}(\tilde{W},\partial\tilde{W})}$}{{...}}
[…]
Remark 3.20.
Note that in the above proof, if $\pi_{1}(M)=\pi_{1}(W)$ is finite, we have $H_{k}(\tilde{W},\partial\tilde{W})\cong H_{k}(D^{2n},S^{2n-1})$ and $\pi_{1}(\tilde{W})=0$ implies by h-cobordism theorem that $\tilde{W}$ is diffeomorphic to a ball [KO24] and $\partial\tilde{W}$ is homeomorphic to a sphere. Therefore such dynamically contact manifolds are necessarily covered by spheres which are the only examples known to the author.
@@ 4. Unit Cotangent bundles and dynamical convexity
If the unit cotangent bundle $ST^{*}M$ of a simply connected manifold $M$ admits a non-degenerate strongly dynamically convex contact form $\lambda$, then the map $\iota_{*}:H_{k}(M;\mathbb{Z}_{2})\rightarrow H_{k}(\mathcal{L}M;\mathbb{Z}_{2})$ induced by the inclusion of constant loops, is an isomorphism for $k\leq n-1$.
@@
Theorem 4.6 ([SER51][LER50]).
@@
Corollary 4.7 ([MCC01]Example 55D).
@@
Since the edge homomorphism $H_{2}(\mathcal{L}M)\rightarrow E^{\infty}_{(2,0)}\rightarrow\cdots\rightarrow E^{2}_{(2,0)}=H_{2}(M)$ in the spectral sequence for the loop-loop fibration coincides with the fibration map $\pi_{*}$ [MIL21, Proposition 64.2] , we have that $E^{2}_{(2,0)}=H_{2}(M)$ persists till the limiting page. This implies that $d_{2}=0$. The fact that the differential $d_{2}$ vanishes forces $E^{2}_{(0,1)}=H_{1}(\Omega M)$ to persist until the last page. The isomorphism $H_{2}(\mathcal{L}M)=H_{2}(M)$ implies that in the limiting page $H_{2}(M)$ is the sole contributor to $H_{2}(\mathcal{L}M)$. Therefore, $H_{1}(\Omega M)=0$.
@@
Theorem 4.10.
The unit cotangent bundle $ST^{*}M$ of a simply connected closed manifold $M$ (not necessarily orientable) with standard contact structure cannot be strongly dynamically convex.
@@
Theorem 4.1 implies $M$ is simply connected, we obtain the result, using theorem 4.10. ∎
Corollary 4.13 (Theorem 1.3).
The unit cotangent bundle $ST^{*}M$ of a closed manifold $M$ cannot be dynamically convex if it is simply connected.
@@ 5. Dynamical convexity and surgery
The idea to obstruct dynamical convexity after surgery on Weinstein fillable contact manifolds is to observe the changes in $SH^{+}$ and in the homology after performing surgery. We refer to [FAU20] for details on how attaching a handle induces maps on the canonical long exact sequence of symplectic homology. In particular, the author shows that, after handle attachment, the Viterbo transfer map 3.11 still respects the action filtration and therefore induces a map on the long exact sequence 3.5.
Note that $SH=\bigoplus_{\gamma\in\pi_{1}}SH_{\gamma}$ where $SH_{\gamma}$ denotes the subgroup generated by orbits in the homotopy class of $\gamma$. In this section we only consider the subgroup generated by contractible orbits and denote it by $SH$ or $SH^{+}$ for the sake of brevity. We consider $SH,SH^{+}$ and homology/cohomology groups over $\mathbb{Q}$ coefficients but omit the coefficients from the notation.
@@
The following proofs in this section also hold for non-Weinstein exact fillable contact manifolds if the fillings have vanishing rational first chern class.
@@
The maps induced by surgery on the canonical long exact sequence [FAU20, section 2.8] are
@@
we note that, since attaching a 1-handle is subcritical, the second and last arrows on symplectic homology are isomorphisms. Also note that since connectedness is preserved $H^{0}$ remains invariant and therefore $\operatorname{rk}(H^{1}(W^{\prime}))=\operatorname{rk}(H^{1}(W))+1$. The map $H^{1}(W^{\prime})\rightarrow H^{1}(W)$, induced by inclusion, is surjective by the long exact sequence in cohomology in section 5. Therefore, by the 4-lemma ([MAC12, Chapter 12 Lemma 3.1]), the map $SH^{+}_{n}(W^{\prime})\rightarrow SH^{+}_{n}(W)$ is a surjection. The vanishing of $SH^{+}_{n}(W^{\prime})$ implies that $SH^{+}_{n}(W)$ vanishes.
@@
If $M$ is simply connected, dynamically convex and of dimension at least $5$, then by [KO24] $M$ is diffeomorphic to a sphere if the filling has vanishing SHSH. This follows from the h-cobordism theorem. Note that vanishing of $SH^{+}$ is independent of the filling by 3.10. ∎
@@
We consider the process of connected sum as attaching a one handle to two components of $M$, thus decreasing the rank of $H^{0}$ by 1 and preserving the ranks of all the higher cohomology groups. $SH$ remains invariant as we are performing subcritical surgery.
Now, if $SH^{+}_{k}(W)=0$ for $k<n+1$, then $SH^{+}_{k}(DT^{*}N)=0$ for $k<n+1$. The proof of lemma 4.2 holds and this implies $H_{k}(N,\mathbb{Z}/2\mathbb{Z})\hookrightarrow H_{k}(\mathcal{L}N,\mathbb{Z}/2\mathbb{Z})$ under inclusion is an isomorphism for $k<n+1$. Since $N$ is simply connected the argument following Lemma 4.2 in section 4 using the Serre spectral sequence would give a contradiction. Therefore there exists $k<n+1$ such that $SH^{+}_{k}(W)=0$.